OpenStudy (anonymous):

Calculate the change in pH when 5.00 mL of 0.100 M HCl(aq) is added to 100.0 mL of a buffer solution that is 0.100 M in NH3(aq) and 0.100 M in NH4Cl(aq).

OpenStudy (mos1635):

initial pH henderson hasselbach formula pH=pKa + log [base]/[acid] [base]=[NH3] [acid]=[NH4Cl] pKa=pKw-pKb (Kw and Kb for water and NH3 are required)

OpenStudy (mos1635):

mole HCl: n1=C1*V1=0.1*0.005=0.0005 moles mole NH3:n2=C2*V2=0.1*0.1=0.01 moles mole NH4Cl: n3=C3*V3=0.1*0.1=0.01 moles HCl + NH3-> NH4Cl initial n1 n2 n3 reacts/produses -n1 -n1 +n1 final 0 n2-n1 n3+n1 final [NH3]=0.0095/0.105 [NH4Cl]= 0.0105/0.105 final pH pH=pKa + log [NH3]/[NH4Cl] so pH1=pKa pH2=pKa + log 0.0095/0.105 ΔpH=log0.0095/0.105 ΔpH=- 1.043

OpenStudy (mos1635):

please check my calc i was not theral

OpenStudy (vincent-lyon.fr):

There is a '0' in pH2 which should not be there. Using millimoles instead of moles avoids working with too many 0's. ΔpH= lg[9.5/10.5] = -0.04 which prooves the buffer effect of this solution.

OpenStudy (mos1635):

@Vincent-Lyon.Fr your resalt makes more sence. althow i am looking to find the extra 0 i don't see it. not in a moode to do it from skrats. can you poit it out for me?

OpenStudy (vincent-lyon.fr):

Sorry, it was not an extra 0, but a missing one. This was ok : final [NH3]=0.0095/0.105 [NH4Cl]= 0.0105/0.105 then pH2 should have read: pH2=pKa + log 0.0095/0.0105 <--- here was the missing 0 I was confused because new volume 0.105 L and amount of NH4+ 0.0105 mol look very much the same.

OpenStudy (mos1635):

yeap got it thanks man